OpenStudy (anonymous):

A current I=2.82cos377t (I in amps, t in seconds, and the "angle" is in radians) flows in a series LR circuit in which L = 3.85 mH and R = 1.16kΩ . What is the average power dissipation in kW?

OpenStudy (anonymous):

I thought 377 isomega so X_L is omega*L=.231

OpenStudy (ybarrap):

Only the resistor dissipates power

OpenStudy (ybarrap):

$$ <I^2\times R> $$

OpenStudy (anonymous):

I tried that and 9.22kW is not the answer

OpenStudy (ybarrap):

Not \(I^2\times R\), but the time average power \(<I^2\times R>\): http://www.wolframalpha.com/input/?i=average+power&lk=4&num=1

OpenStudy (ybarrap):

What's the average of the cos function? Multiply that times 9.22 kW

OpenStudy (ybarrap):

Divide by 2pi, a typical period to find average

OpenStudy (anonymous):

How do you get the average of cos?

OpenStudy (ybarrap):

I meant \(\cos^2x\): $$ <I^2\times R>=\cfrac{1}{2\pi}\int_{-\pi}^{\pi}(2.82\cos(377~t))^2\times 1160~dt $$

OpenStudy (ybarrap):

$$ \int_{-\pi}^{\pi}\cos^2(377~t)~dt=\pi $$

OpenStudy (ybarrap):

$$ <I^2\times R>=\cfrac{2.82^2\times 1,160}{2\pi}\times \pi\\ =\cfrac{2.82^2\times 1,160}{2}\\ $$

OpenStudy (anonymous):

Oh ok I see, thank you! I have been trying this one for an hour!

OpenStudy (ybarrap):

So, because we have a sinusoidal function, the dissipated power is actually 1/2 of that of a DC circuit, which makes sense that its less, but that it's 1/2 is surprising.

OpenStudy (ybarrap):

Oh, btw, that is why dimmer switches use variable inductors and not variable resistors -- less power dissipated.

OpenStudy (anonymous):

Thank you :)